You’ve observed the following returns on Crash-n-Burn Computer’s stock over the past five years: 8 percent, −15 percent, 19 percent, 31 percent, and 21 percent. Suppose the average inflation rate over this period was 3.1 percent and the average T-bill rate over the period was 3.9 percent.

Every time I look this up, I can only find the real return, not the real risk free return....

a.

What was the average real risk free rate over this time period? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. What was the average real risk premium? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a. Average real risk free rate =(1+3.9%)/(1+3.1%)-1 =0.78%

b.Average Return =(8%-15%+19%+31%+21%)/5 =12.80%
Average Real risk free rate =(1+12.80%)/(1+3.1%)-1 =9.4083%
Average real risk premium =9.4083%-0.78% =8.63%